Week of 170904

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Vladimir (LHCb), Marian (Network), Julia (WLCG), Xavier (chair, Storage), Maarten (Alice, WLCG), Alberto (Monitoring), Andrea (MW, Storage), Nils (Computing), Andrei (DB), Vincent (Security)
  • remote: Luca (CNAF), Nurcan (ATLAS), Sang Un (KISTI), John (RAL), David (IN2P3), Christoph (CMS), David (FNAL), Dmitry (KIT), Vincenzo (EGI)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Production is running with ~320k slots of running jobs, with NERSC_Cori HPC providing good number of slots as well.
    • bigpanda.cern.ch monitor has been migrated to https from http today (w/o CERN SSO login requirement, the issue is being discussed in INC:1446438 "Chrome not supported for accessing BigPanDA using SSO").
    • The service monitoring is not working, reported in GGUS:130354, announcement from IT-monitoring team: Data is arriving with some delay into Meter and Timber service.
  • CMS reports ( raw view) -
    • Situation very much like last week
      • High CPU utilization (~120-140k cores over last few days)
      • Much activity in the data transfer system (incl. staging from tape)
    • Issues being followed with Tier-0 and Tier-1 admins
      • SRM of CMS Castor instance seems to have an issue (time outs): GGUS:130363
        • CMS should probably check and reduce its SRM dependence...
      • Transfer/staging issues at CNAF
      • Transfer issues at CCIN2P3
        • Low transfer rates (GGUS:130332 - suffering from bug in FTS 3.5 at RAL, perhaps more reasons)
      • Heavy staging at RAL (GGUS:130193)
      • Transfer timeouts at FNAL (GGUS:130234)
      • Low transfer rates from CERN to KIT (GGUS:130160 - Can possibly be improved by changes to local Phedex agents, any updates?)

  • ALICE -
    • Continued very high activity
    • CERN: team ticket GGUS:130315 opened Thu
      • 3 CREAM CEs had their BDII stuck
      • It made ALICE VOBOXes stop submitting jobs

  • LHCb reports ( raw view) -
    • Activity
      • Monte Carlo simulation, data processing and user analysis
    • Site Issues
      • T0:
        • Incomplete python installation at worker nodes (GGUS:130018)
      • T1:
        • Failed transfers from IC to SARA (IPV6) (GGUS:129946); no news
        • Failed transfers from many sites to dCache sites, see (GGUS:130190); Resolved by using proper parameter in SRM
        • We have peak of failed transfers at EOS every day at 5:00 (GGUS:130335)

Sites / Services round table:

  • ASGC:
  • BNL:
  • CNAF: NTR
  • EGI: NTR
  • FNAL: Observed syncing issues between OSG-SNOW and GGUS when a ticket is re-opened (investigating)
  • IN2P3: Investigating issue reported by CMS, some PhEDEx parameters have been changed with no significant results. As suggested by CMS in GGUS ticket, we also tried to change from using FTS@RAL to FTS@CERN, but this fails.
  • JINR: All looks normal, but we see somewhat low pilot efficiency (sometimes only 1 task on 10 reserved cores).
  • KISTI: NTR
  • KIT: NTR
  • NDGF:
  • NL-T1: Due to a scheduling conflict we are unable to connect. Currently a part of our files on tape is unavailable because of a robotic failure that started this weekend. We hope to solve this today.
  • NRC-KI:
  • OSG:
  • PIC: NTR
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: NTR (apart from possible fallout from network/ceph incident today)
  • CERN storage services:
    • FTS: FTS prod cluster upgraded today to v3.7.3.
  • CERN databases:
    • CMSONR ADG in Wigner was down of Friday due to an issue with storage
    • Standby database for ATLAS offline ADCR_ADG lagging behind production OTG:0039485
    • Some Oracle development and integration database intermittently unavailable OTG:0039558
  • GGUS: NTR
  • Monitoring:
    • Draft Site Availability report for AUG 2017 distributed.
    • Following OTG:0039584
  • MW Officer: NTR
  • Networks:
    • Oxford - GGUS:130032 - Further investigation has shown that there was a network issue at Oxford during August, but performance has improved recently.
    • SARA/IC - GGUS:129946 - Load balancing issue, waiting on EduPERT to respond
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r22 < r21 < r20 < r19 < r18 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r22 - 2017-09-11 - MarianBabik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback